# Plastic tensor networks for interpretable generative modeling

### v1

@@ 4 Summary and Discussion
In order to obtain interpretable tensor trees for classical distributions, we replaced the double layer in the previous BMATT scheme [11] by a single layer with nonnegative tensors. The proposed NATT scheme uses nonnegative projected gradient descent and NMF to respect a nonnegative constraint on the variational parameters. We demonstrated that the new scheme performs almost as well as the double-layer scheme in minimizing the NLL and identifying the relational structures. By using the double-layer scheme to provide the single-layer scheme with an initial network structure, we can more consistently obtain good candidate structures than when considering only the single-layer scheme alone, while still keeping the network interpretable as a probabilistic graph model. We also showed that the NATT and BMATT schemes are able to solve a variety of learning problems, including the XOR problem, which is characterized by an absence of two-point correlations in the input data, a random Bayesian network learning problem, where only a subset of sites are visible and the internal structure of the Bayesian network must be identified, and a real-world problem in phylogenetics, where mtDNA sequences are used to uncover a phylogenetic tree linking different members of order Carnivora. While the main advantage of the BMATT scheme centers around generative model quality (and by extension, sample quality) as measured by the NLL, the NATT scheme we propose in this work is useful when a probabilistic model is desirable: even if the BMATT scheme gives the correlation structure, it cannot provide an interpretable model.
@@
Finding the best structure to represent the data also provides clustering information in the process. Knowing the joint distribution $P(x)$ describing the data and being able to easily trace out variables (i.e. when any subgraph of the network representing $P(x)$ can be contracted efficiently, which is true when $P(x)$ is a tree) gives complete knowledge of the correlation structure of the data. For complicated data, we can leverage the BMATT scheme to find a good candidate initial structure and then use the NATT scheme to provide a compact representation of $P(x)$ that is entirely explainable classically. In this manner, we can benefit from the strengths of both schemes.
@@
5 Acknowledgments
This work was partially supported by the joint project of Kyoto University and Toyota Motor Corporation, titled “Advanced Mathematical Science for Mobility Society.” K. A. would like to acknowledge the support of the Global Science Graduate Course (GSGC) program of the University of Tokyo. K. H. acknowledges the support from JSPS KAKENHI (Grant No. 20K03766 and 24K06886) and a Grant-in-Aid for Transformative Research Areas “The Natural Laws of Extreme Universe—A New Paradigm for Spacetime and Matter from Quantum Information” (KAKENHI Grants No. 21H05182 and No. 21H05191) from JSPS of Japan. T. O. acknowledges the support from JSPS KAKENHI (Grant No. 23H03818 and 22K18682), the Endowed Project for Quantum Software Research and Education, the University of Tokyo (https://qsw.phys.s.u-tokyo.ac.jp/), and The Center of Innovations for Sustainable Quantum AI (JST Grant No. JPMJPF2221). N. K. acknowledges the support from JSPS KAKENHI (Grant No. 23H01092). Some of the computation in this work have been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.

### v2

@@ IV Summary and Discussion
In order to obtain interpretable tensor trees for classical distributions, we replaced the double layer in the previous BMATT scheme [14] by a single layer with nonnegative tensors. The proposed NATT scheme uses nonnegative projected gradient descent and NMF to respect a nonnegative constraint on the variational parameters. We demonstrated that the new scheme performs almost as well as the double-layer scheme in minimizing the NLL and identifying the relational structures. By using the double-layer scheme to provide the single-layer scheme with an initial network structure, we can more consistently obtain good candidate structures than when considering only the single-layer scheme alone, while still keeping the network interpretable as a probabilistic graphical model. We also showed that the NATT and BMATT schemes are able to solve a variety of learning problems, including the XOR problem, which is characterized by an absence of two-point correlations in the input data, a random Bayesian network learning problem, where only a subset of sites are visible and the internal structure of the Bayesian network must be identified, and a real-world problem in phylogenetics, where mtDNA sequences are used to uncover a phylogenetic tree linking different members of order Carnivora. While the main advantage of the BMATT scheme centers around generative model quality (and by extension, sample quality) as measured by the NLL, the NATT scheme we propose in this work is useful when a probabilistic model is desirable: even if the BMATT scheme gives the correlation structure, it cannot provide an interpretable model.
@@
Finding the best structure to represent the data also provides clustering information in the process. Knowing the joint distribution $P(x)$ describing the data and being able to easily trace out variables (i.e. when any subgraph of the network representing $P(x)$ can be contracted efficiently, which is true when $P(x)$ is a tree) gives complete knowledge of the correlation structure of the data. For complicated data, we can leverage the BMATT scheme to find a good candidate initial structure and then use the NATT scheme to provide a compact representation of $P(x)$ that is entirely explainable classically. In this manner, we can benefit from the strengths of both schemes. We note that with datasets with a composition that changes over time, these plastic ATT approaches are able to adapt in accordance with the input data.
@@
V Data Availability Statement
The data and code used to run the numerical experiments presented in this work can be found at the following URL: https://github.com/redpinetree/att-ml.
Acknowledgements.
This work was partially supported by the joint project of Kyoto University and Toyota Motor Corporation, titled “Advanced Mathematical Science for Mobility Society.” K. A. would like to acknowledge the support of the Global Science Graduate Course (GSGC) program of the University of Tokyo. K. H. acknowledges the support from JSPS KAKENHI (Grant No. 20K03766 and 24K06886) and a Grant-in-Aid for Transformative Research Areas “The Natural Laws of Extreme Universe—A New Paradigm for Spacetime and Matter from Quantum Information” (KAKENHI Grants No. 21H05182 and No. 21H05191) from JSPS of Japan. T. O. acknowledges the support from JSPS KAKENHI (Grant No. 23H03818 and 22K18682), the Endowed Project for Quantum Software Research and Education, the University of Tokyo (https://qsw.phys.s.u-tokyo.ac.jp/), and The Center of Innovations for Sustainable Quantum AI (JST Grant No. JPMJPF2221). N. K. acknowledges the support from JSPS KAKENHI (Grant No. 23H01092). N.K. also appreciates useful comments from H. Shinaoka. Some of the computation in this work have been done using the facilities of the Supercomputer Center, the Institute for Solid State Physics, the University of Tokyo.